# DP-BREM: Differentially-Private and Byzantine-Robust Federated Learning with Client Momentum

### v2

@@ 5.2 Design of DP-BREM+
[…]
Figure 3: MNIST - Varying the percentage of Byzantine clients $\delta_{B}$ (with $\epsilon=3$).
Figure 4: CIFAR-10 - Varying the percentage of Byzantine clients $\delta_{B}$ (with $\epsilon=4$).
@@ 5.3 Security Analysis
where $\{z_{i}\}_{i\in\mathcal{I}_{H}}$ denotes the input of all the honest clients, $\mathsf{Real}$ denotes a random variable representing the joint view of all the parties in the protocol’s execution, $\Omega_{\mathcal{I}_{M}\cup\mathsf{S}}$ indicates a polynomial-time algorithm implementing the "next-message" function of the parties in $\mathcal{I}_{M}\cup\mathsf{S}$ (see [35, Appendix 11.5]), and $\equiv_{\mathsf{C}}$ denotes computational indistinguishability. In summary, the server and clients do not learn anything besides the final aggregated result.
@@ 6 Experimental Evaluation
Figure 5: MNIST - Varying privacy budget $\epsilon$ (with $\delta_{B}=30\%$ Byzantine clients).
Figure 6: CIFAR-10 - Varying privacy budget $\epsilon$ (with $\delta_{B}=15\%$ Byzantine clients).
In this section, we demonstrate the effectiveness of the proposed DP-BREM/DP-BREM+ on achieving both good privacy-utility tradeoff and Byzantine robustness via experimental results on MNIST [25] and CIFAR-10 [24] datasets with non-IID setting (refer to Appendix J.1 for more details on the datasets and model architectures). All experiments are developed via PyTorch [^1] Our source code will be available after the acceptance of the paper..
Byzantine Attacks. We consider four existing Byzantine attacks in our experiments, including ALIE ("a little is enough") [3], IPM (inner-product manipulation) [45], LF (label-flipping), and the state-of-the-art MTB ("manipulating-the-Byzantine") [38]. Refer to Appendix J.1 for more details.
Compared Methods. We compare the performance of six approaches against Byzantine attacks, including DP-BREM/+ (our approach) [^2] Since DP-BREM+ achieves the same DP and robustness guarantees as DP-BREM, we did not perform the empirical experiments with secure aggregation because the accuracy results will be exactly the same as DP-BREM. We use DP-BREM/+ to denote both DP-BREM and DP-BREM+, and the implementation follows Algorithm 1., a variant of DP-FedSGD [30] with both record and client norm clipping, DDP-RP [43], DP-RSA [50], a variant of CM [47] with DP noise, and DP-LFH. The comparison (on trust assumption and mechanism overview) of these approaches is provided in Table 1, and Appendix J.1 shows more details of each approach. In summary, DP-BREM/+, DP-FedSGD, and DP-CM add central noise to the aggregation, but DP-BREM+ does not require a trusted server due to the secure aggregation technique. DDP-RP adds partial local noise to the client’s update with secure aggregation. DP-RSA and DP-LFH add local noise to the client’s update. We fix $\delta=10^{-6}$ for $(\epsilon,\delta)$-DP in all experiments. For the setting of other parameters, refer to Appendix J.2.
@@ 6.1 Robustness Evaluation with DP
We consider a fixed privacy budget $\epsilon$ and implement each attack with different percentages of Byzantine clients $\delta_{B}=\frac{|\mathcal{B}|}{n}$ for the four attacks, and compare the accuracy among all approaches. The results for MNIST (with $\epsilon=3$) and CIFAR-10 (with $\epsilon=4$) datasets are shown in Figures 3 and 4. Though the detailed results differ under different attacks and for two datasets, we have some general observations:
@@
4) CM with DP noise (or DP-CM) has a relatively small accuracy decrease for a relative small $\delta_{B}$. It is the benefit of the median-based robust aggregator. But the sensitivity is larger than the average-based aggregators, as discussed in Example 1, the aggregated DP noise is too large to obtain a high accuracy, even when $\delta_{B}=0$.
@@
6) DP-RSA has relatively poor accuracy compared with other approaches, even when $\delta_{B}=0$. It is caused by the sign-SGD aggregator, which only aggregates element-wise signs instead of values, leading to information loss. Moreover, the local DP noise makes the Byzantine attacks easier to succeed.
6.2 Privacy-Utility Tradeoff with Attack
@@
We consider a fixed percentage of Byzantine clients $\delta_{B}$ for each attack under different values of privacy budget $\epsilon$, and compare the accuracy of all approaches. The results for MNIST (with $\delta_{B}=30\%$) and CIFAR-10 (with $\delta_{B}=15\%$) datasets are shown in Figures 5 and 6. For both datasets, we consider four different levels of privacy, where ϵ=infitalic-$\epsilon=\inf$ means the standard deviation of DP noise is 0, but we still implement record-level clipping to illustrate how the noise affects the results while keeping other settings including the clipping step the same. Though the detailed results differ under different attacks and for two datasets, DP-BREM/+ has the best accuracy among all approaches, especially under IPM and MTB attacks. Note that when $\sigma=0$ for DP noise (i.e., ϵ=infitalic-$\epsilon=\inf$), both DP-BREM/+ and DP-LFH reduce to LFH, thus they have the same results in this case. We can observe that with a moderate privacy budget, such as $\epsilon\geqslant 2$, DP noise only has a negligible impact on the accuracy. But if $\epsilon$ is too small, such as $\epsilon=1$ in Figure 5, DP-BREM/+ suffers a relatively larger impact (but still acceptable) from DP noise. Note that when there exist Byzantine attacks, reducing the DP noise to $\sigma=0$ (i.e., ϵ=infitalic-$\epsilon=\inf$) does not significantly improve the accuracy of DP-BREM/+ compared with ϵ<infitalic-$\epsilon<\inf$, because the performance is largely impacted by Byzantine clients’ perturbations. However, the accuracy of DP-LFH is greatly reduced when ϵ<infitalic-$\epsilon<\inf$, since the local DP noise impacts the robustness of the aggregator. This observation is consistent with our theoretical analysis in Limitation 2 of DP-LFH (Section 3.2).
6.3 Other Results
@@
Table 3: Running time1 (in milliseconds) per round per client in MNIST dataset
@@
Impact of $R$ in DP-BREM/+. Figure 7 (in Appendix J) shows how the accuracy changes w.r.t. the record-level clipping bound $R$ in DP-BREM/+. The results demonstrate that when there are fewer Byzantine clients (i.e., smaller $\delta_{B}$) or the noise multiplier $\sigma$ is smaller (i.e., larger $\epsilon$), we need to set a larger $R$ to obtain better accuracy. This observation is consistent with the theoretical analysis of parameter tuning discussed in Theorem 2 and its interpretation.
Impact of $q$ in DP-BREM/+. In all previous experiments, we set client-level sampling rate $q=1$ by default. As discussed in Sec. 4.1, aggregating a subset $I_{t}$ of clients in (12) is one of the major differences from LFH. In Table 4, we show how the utility can be improved by this design under different attack percentages $\delta_{B}$ if we can choose an optimal $q$ (which can be regarded as a tunable parameter). Intuitively, when there is no attack, a smaller $q$ can provide more privacy amplification, i.e., a smaller $\sigma$ is needed for the same value of $\epsilon$ in DP; but if $q$ is too small, the small aggregate population will lead to a larger variance of the aggregation. When there exists Byzantine attacks, a smaller value of $q$ can reduce the attack impact for each round because only a portion of Byzantine clients are selected for aggregation. Therefore, with increased $\delta_{B}$, the optimal $q$ (highlighted in Table 4) is decreased.
Table 4: Model accuracy when varying $q$ of DP-BREM/+ with $\epsilon=2$ under MTB attack in CIFAR-10 dataset.
@@ 7 Related Work
Due to limited space, we only discuss the most relevant defenses below and put other related work in Appendix K. Other works either only achieve DP or Byzantine robustness (but not both), or combine secure aggregation with Byzantine robustness without realizing DP.
Wang et al. [43] proposed an FL scheme (DDP-RP) to provide Distributed DP (via encryption) and robustness (via range proof technologies); however, this scheme only verifies whether the local model weights are in a bounded range, which provides weak robustness. In comparison, our solution utilizes client momentum and centered clipping to guarantee Byzantine robustness with provable convergence analysis. Zhu et al. [50] replaces the value aggregation with sign aggregation, which provides robustness because each client has a limited impact on the aggregation. The DP noise is added to the local gradient before the sign operation. Since it only aggregates the element-wise sign (instead of the value) of clients’ gradients, it has degraded convergence due to information loss. Also, [50] only accounts for the privacy cost of one iteration, instead of the composition of all iterations in FL. Thus, the privacy cost computed in [50] is underestimated. As a comparison, our solution is based on the original SGD (with momentum), and we account for the privacy cost of all iterations. Our experimental results have confirmed that DP-BREM outperforms both of these approaches.
@@ 8 Conclusions
Appendix A Proof of Lemma 2 (Aggregation Sensitivity)

### v3

@@ 5.2 Design of DP-BREM+
Figure 3: With fixed privacy budget $\epsilon$, varying the percentage of Byzantine clients $\delta_{B}$ for three datasets.
@@ 5.3 Security Analysis
where $\{z_{i}\}_{i\in\mathcal{I}_{H}}$ denotes the input of all the honest clients, $\mathsf{Real}$ denotes a random variable representing the joint view of all the parties in the protocol’s execution, $\Omega_{\mathcal{I}_{M}\cup\mathsf{S}}$ indicates a polynomial-time algorithm implementing the "next-message" function of the parties in $\mathcal{I}_{M}\cup\mathsf{S}$ (see [36, Appendix 11.5]), and $\equiv_{\mathsf{C}}$ denotes computational indistinguishability. In summary, the server and clients learn nothing besides the final aggregated result.
@@ 6 Experimental Evaluation
Figure 4: With fixed percentage of Byzantine clients $\delta_{B}$, varying privacy budget $\epsilon$ for three datasets.
In this section, we demonstrate the effectiveness of the proposed DP-BREM/DP-BREM+ on achieving both good privacy-utility tradeoff and Byzantine robustness via experimental results on MNIST [26], CIFAR-10 [25], and FEMNIST [8] datasets with non-IID setting (refer to Appendix E for more details on the datasets and model architectures). Note that MNIST and CIFAR-10 have 10 classes, while FEMNIST includes 62 classes. All experiments are developed via PyTorch [^1] Our source code is available at https://github.com/xiaolangu/DP-BREM.
Byzantine Attacks. We consider four existing Byzantine attacks in our experiments, including ALIE ("a little is enough") [3], IPM (inner-product manipulation) [46], LF (label-flipping), and the state-of-the-art MTB ("manipulating-the-Byzantine") [39]. Refer to Appendix E for more details.
Compared Methods. We compare the performance of six approaches against Byzantine attacks, including DP-BREM/+ (our approach) [^2] Since DP-BREM+ achieves the same DP and robustness guarantees as DP-BREM, we did not perform the empirical experiments with secure aggregation because the accuracy results will be exactly the same as DP-BREM. We use DP-BREM/+ to denote both DP-BREM and DP-BREM+, and the implementation follows Algorithm 1., a variant of DP-FedSGD [31] with both record and client norm clipping, DDP-RP [44], DP-RSA [51], a variant of CM [48] with DP noise, and DP-LFH. The comparison (on trust assumption and mechanism overview) of these approaches is provided in Table 1, and Appendix E shows more details of each approach. In summary, DP-BREM/+, DP-FedSGD, and DP-CM add central noise to the aggregation, but DP-BREM+ does not require a trusted server due to the secure aggregation technique. DDP-RP adds partial local noise to the client’s update with secure aggregation. DP-RSA and DP-LFH add local noise to the client’s update. We fix $\delta=10^{-6}$ for $(\epsilon,\delta)$-DP in all experiments. For the setting of other parameters, refer to Appendix F.
@@ 6.1 Robustness Evaluation with DP
We consider a fixed privacy budget $\epsilon$ and implement each attack with different percentages of Byzantine clients $\delta_{B}=\frac{|\mathcal{B}|}{n}$ for the four attacks, and compare the accuracy among all approaches. We note that a complex dataset requires a more sophisticated model architecture and makes it more challenging to maintain good utility in the presence of DP and Byzantine attacks. Therefore, in our experiments with CIFAR-10 (which has three color channels) and FEMNIST (which includes 62 classes), we use slightly larger $\epsilon$ values and a smaller number of Byzantine clients. These choices are still within a reasonable range. Previous papers, such as [1] and [50], also used larger privacy budgets for the CIFAR-10 dataset compared to the MNIST dataset. The results for MNIST (with $\epsilon=3$), CIFAR-10 (with $\epsilon=4$), and FEMNIST (with $\epsilon=4$) datasets are shown in Figure 3. Compared to the results on the MNIST dataset (with 10 classes), the accuracy on the FEMNIST dataset is lower due to the larger number of classes. Though the detailed results vary under different attacks and across three datasets, we have some general observations:
@@
4) CM with DP noise (or DP-CM) has a relatively small accuracy decrease for a relatively small $\delta_{B}$. It is the benefit of the median-based robust aggregator. But the sensitivity is larger than the average-based aggregators, as discussed in Example 1, the aggregated DP noise is too large to obtain a high accuracy, even when $\delta_{B}=0$.
@@
6) DP-RSA has relatively poor accuracy compared to other approaches, even when $\delta_{B}=0$. This is caused by the sign-SGD aggregator, which only aggregates element-wise signs rather than the full precision gradient, leading to large information loss. Moreover, the local DP noise makes Byzantine attacks easier to succeed. We note that DP-RSA does not converge for the FEMNIST dataset (possibly caused by the sign aggregation), even without DP noise and Byzantine clients, and thus we do not present the results for this dataset.
7) Under the ALIE attack, it is possible for a small number of Byzantine clients to improve the accuracy of the model compared to the scenario without any Byzantine clients. For instance, an ALIE attack with $10\%$ Byzantine clients can achieve higher accuracy than the case with $0\%$ Byzantine clients across all defense aggregators except DP-LFH. This improvement occurs because the ALIE attack generates malicious gradients that are close to the averaged good gradients but deviate slightly using a scaling factor. This factor is determined based on the total number of clients and the proportion of Byzantine clients, designed to bypass any anomaly detection mechanisms employed by the central server. Consequently, when the number of Byzantine clients is relatively small, the malicious gradients can enhance model accuracy compared to benign gradients, where the record-level clipping is used to achieve DP.
6.2 Privacy-Utility Tradeoff under Attack
@@
We consider a fixed percentage of Byzantine clients $\delta_{B}$ for each attack under different values of privacy budget $\epsilon$, and compare the accuracy of all approaches. The results for MNIST (with $\delta_{B}=30\%$), CIFAR-10 (with $\delta_{B}=15\%$), and FEMNIST (with $\delta_{B}=15\%$) datasets are shown in Figure 4. For all three datasets, we consider four different levels of privacy, where ϵ=infitalic-$\epsilon=\inf$ means the standard deviation of DP noise is 0. However, we still implement record-level clipping to illustrate how the noise affects the results while keeping other settings including the clipping step the same.
It’s essential to highlight that while the privacy-utility curve is generally monotonic in the absence of Byzantine attacks, this may not hold under Byzantine attacks due to two sources of perturbation. When malicious perturbation dominates, the impact of DP noise on utility is typically minimal. Additionally, different defense aggregators exhibit varying sensitivities to malicious perturbation and DP noise across various datasets, even when the number of malicious clients and $\epsilon$ values are the same. Consequently, observations can vary across defense aggregators, attacks, and datasets (with different parameters). For example, DP noise has a very small (or almost negligible) impact on DP-FedSGD compared to DP-BREM. This could be because DP-FedSGD aggregates more information than the momentum-based solution, leading to a better signal-to-noise ratio (SNR) and thus greater robustness to DP noise. However, the attack has a more significant impact on DP-FedSGD.
Though the detailed results vary under different attacks and across the three datasets, DP-BREM/+ generally achieves the highest accuracy among almost all approaches, especially under IPM and MTB attacks. The only exception is when $\epsilon=2$ for the FEMNIST dataset, where the accuracy of DP-BREM/+ is lower than that of DP-FedSGD. This is because the client momentum in DP-BREM/+ restricts the information that can be learned from each new iteration, making the increased DP noise have a greater impact on the model’s accuracy.
Note that when $\sigma=0$ (i.e., ϵ=infitalic-$\epsilon=\inf$), both DP-BREM/+ and DP-LFH reduce to LFH, thus they have the same results in this case. We can observe that with a moderate privacy budget, such as $\epsilon\geqslant 2$, DP noise only has a negligible impact on the accuracy. But if $\epsilon$ is too small, such as $\epsilon=1$ for the MNIST dataset in Figure 4, DP-BREM/+ suffers a relatively larger impact (but still acceptable) from DP noise. Note that when there exist Byzantine attacks, reducing the DP noise to $\sigma=0$ (i.e., ϵ=infitalic-$\epsilon=\inf$) does not significantly improve the accuracy of DP-BREM/+ compared with ϵ<infitalic-$\epsilon<\inf$, because Byzantine clients’ perturbations largely impact the performance. However, the accuracy of DP-LFH is greatly reduced when ϵ<infitalic-$\epsilon<\inf$, since the local DP noise impacts the robustness of the aggregator. This observation is consistent with our theoretical analysis in Limitation 2 of DP-LFH (Section 3.2).
@@
Table 3: Running time1 (in milliseconds) per round per client on the MNIST dataset.
@@
6.3 Other Results
Table 4: Model accuracy when varying $C$ of DP-BREM/+ with $\epsilon=4$ under IPM and MTB attacks on FEMNIST dataset.
$\delta_{B}$ $C=0.5$ $C=1$ $C=2$ $C=3$ $C=4$
0% 0.622 0.647 0.625 0.621 0.627
IPM 10% 0.407 0.524 0.555 0.528 0.514
IPM 20% 0.060 0.305 0.436 0.413 0.392
MTB 10% 0.591 0.605 0.535 0.525 0.545
MTB 20% 0.554 0.537 0.477 0.426 0.426
Impact of $R$ in DP-BREM/+. Figure 5 shows how the accuracy changes w.r.t. the record-level clipping bound $R$ in DP-BREM/+. The results demonstrate that when there are fewer Byzantine clients (i.e., smaller $\delta_{B}$) or the noise multiplier $\sigma$ is smaller (i.e., larger $\epsilon$), we need to set a larger $R$ to obtain better accuracy. This observation is consistent with the theoretical analysis of parameter tuning discussed in Theorem 2 and its interpretation.
Impact of $C$ in DP-BREM/+. We use the fixed client-level clipping bound $C$ for each dataset in previous experiments. Table 4 illustrates how varying values of $C$ (while keeping the default and fixed $R$) can influence model accuracy. In the absence of Byzantine attacks, the value of $C$ has a relatively small impact on the model accuracy. However, in the presence of Byzantine attacks, the effect of $C$ varies depending on the nature of the attack. For instance, attacks like the IPM attack, which deviate significantly from benign gradients, benefit from a slightly larger $C$ as it allows more useful information (from benign clients) to be retained. Conversely, for attacks like the MTB attack, which aim to evade detection by aligning more closely with benign gradients, a slightly smaller $C$ can improve accuracy by reducing the impact of the attack on the aggregated gradient.
Impact of $q$ in DP-BREM/+. In previous experiments, we set client-level sampling rate $q=1$ by default. As discussed in Sec. 4.1, aggregating a subset $I_{t}$ of clients in (12) is one of the major differences from LFH. Table 5 demonstrates the utility improvement by optimizing $q$ under different attack percentages $\delta_{B}$. Intuitively, without attacks, a smaller $q$ enhances privacy amplification, reducing the required $\sigma$ for a given $\epsilon$ in DP; however, too small a $q$ increases aggregation variance. Under Byzantine attacks, a smaller $q$ mitigates attack impact as only a subset of Byzantine clients are aggregated. Thus, with higher $\delta_{B}$ the optimal $q$ (highlighted in Table 5) decreases.
Table 5: Model accuracy when varying $q$ of DP-BREM/+ with $\epsilon=2$ under MTB attack on the CIFAR-10 dataset.
@@ 7 Related Work
Due to limited space, we only discuss the most relevant defenses below and put other related work in Appendix M. Other works either only achieve DP or Byzantine robustness (but not both), or combine secure aggregation with Byzantine robustness without realizing DP.
Wang et al. [44] proposed DDP-RP, an FL scheme offering Distributed DP (via encryption) and robustness (via range-proof technologies). However, this scheme only verifies if local model weights are within a bounded range, providing weak robustness. Our solution, in contrast, employs client momentum and centered clipping for Byzantine robustness with provable convergence. Zhu et al. [51] uses sign aggregation for robustness, thus each client has limited impact, and adds DP noise to local gradients before sign operations. This method suffers from information loss, resulting in degraded convergence, and only accounts for the privacy cost of one iteration, underestimating the overall cost. Our solution, based on original SGD with momentum, considers the privacy cost of all iterations. Experimental results show that DP-BREM outperforms both approaches.
@@ 8 Conclusions
Acknowledgments
The authors would like to thank the anonymous reviewers and the shepherd for their valuable comments and suggestions. Li Xiong was partly supported by NSF grants CNS-2124104, CNS-2125530, IIS-2302968, and NIH grants R01LM013712, R01ES033241.
Appendix A Proof of Theorem 1 (Privacy Analysis)
Proof.
Also, due to the client-level sampling (i.e., each client was selected by the server w.p. $q$) and record-level sampling (i.e., each record was selected by client $\mathsf{C}_{i}$ w.p. $p_{i}$), the overall sampling rate is $qp_{i}$. Then, by applying the privacy accountant of Gaussian DP [12] (Lemma 10 in Appendix I), DP-BREM satisfies $\mu_{i}$-GDP with $\mu_{i}$ shown in (16). Finally, by converting $\mu_{i}$-GDP to $(\epsilon_{i},\delta)$-DP via Lemma 9, we get (15), which finishes the proof. ∎
Appendix B Proof of Theorem 2 (Aggregation Error)
Proof.
Our proof heavily relies on several useful lemmas shown in Appendix G, where Lemma 3 splits the L2-norm of summation of vectors into weighted summation of vectors’L2-norm, and Lemma 4 provides the optimal strategy to choose these weights.
where
=$\overset{(\mathsf{b})}{=}[\rho R/(R+c)+\phi]^{2}$
$\leqslant|\mathcal{H}_{t}|\cdot[\rho R/(R+c)+\phi]^{2}$
where $(\mathsf{a})$ used the assumption that $\{\bar{\bm{m}}_{t,i}\}_{i\in\mathcal{H}_{t}}$ are independent, then the random variables $\{\bm{y}_{i}\}_{i\in\mathcal{H}_{t}}$ are also independent; $(\mathsf{b})$ used contractivity of a clipping (projection) step. Thus,
Putting into Together. Combining all terms, we have
$\qquad\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$
The Final Result. On the other hand, we have
$\qquad\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{m}_{t}^{*}\|^{2}$
where $(\mathsf{a})$ is obtained by using Lemma 3 for any $\gamma\in(0,1)$; $(\mathsf{b})$ is obtained from (B) and (B), where $\Phi$ is defined in (B); $(\mathsf{c})$ is obtained by taking $\gamma=\frac{\Phi}{\Phi+\frac{\rho}{\sqrt{|\mathcal{H}|}}}$. Furthermore, if we assume $\phi\leqslant O(\rho)$ and $\tau\leqslant O(\rho)$, we can rewrite (B) as the following version
which finishes the proof of Theorem 2. ∎
Appendix C Proof of Theorem 3 (Convergence Rate)
Proof.
Comparing with the aggregation error of $O(\rho^{2}|\mathcal{B}|/n)$ (ignoring constants and higher order terms) in [23, Lemma 9], our aggregation error shown in (19) replaces the term $|\mathcal{B}|$ by $|\mathcal{B}|+\sqrt{d}\sigma/q$, which means a slower convergence due to DP noise. Then, following the result in [23, Theorem VI] and its informal version in (5), we get the convergence rate of our algorithm as in (20). Note that our aggregation utilizes a client-level sampling rate $q$, i.e., approximate $nq$ clients participate in the aggregation for one iteration. We need to replace the term of $\frac{1}{n}$ in (5) by $\frac{1}{nq}$ in (20). ∎
Appendix D Detailed Steps of DP-BREM+ in Figure 2
[…]
Appendix E Experimental Setup
Datasets (non-IID) and Model Architecture. We use three datasets for our experiments: MNIST [26] CIFAR-10 [25] and FEMNIST [8], where the number of total clients is $n=100$ for the former two datasets, and $n=400$ for FEMNIST dataset. Note that the MNIST and CIFAR-10 datasets only have 10 classes, while the FEMNIST dataset has 62 classes (including 10 digits, 26 lowercase letters, and 26 uppercase letters). For the MNIST dataset, we use the CNN model from PyTorch example [^4] https://github.com/pytorch/opacus. For the CIFAR-10 dataset, we use the CNN model from the TensorFlow tutorial [^5] https://www.tensorflow.org/tutorials/images/cnn, like the previous works [50, 31]. To simulate the heterogeneous data distributions, we make non-i.i.d. partitions of the datasets, which is a similar setup as [50] and is described below. For the FEMNIST dataset, we use a CNN model with 2 convolution layers and 2 fully connected layers.
3) Non-IID FEMNIST: The FEMNIST dataset is pre-partitioned based on the writer of the characters, simulating a non-IID scenario. Each client’s local dataset consists of samples written by individual users, introducing variability in handwriting styles. We use the TensorFlow-Federated API [^6] https://www.tensorflow.org/federated/api_docs/python/tff/simulation/datasets/emnist/load_data to load the first 400 partitions, representing data from 400 clients. Unlike the MNIST dataset, which includes only digits, FEMNIST includes both digits and uppercase and lowercase letters, spanning 62 classes (10 digits + 52 letters).
3) LF (label-flipping). The attacker modifies the labels of all examples of corrupted clients’ data and computes a new gradient, then uses a gradient replacement strategy (similar to [2]) to enhance the impact on the global model. Specifically, the attacker computes a benign gradient $\mathbf{g}_{\text{benign}}$ with non-flipped labels and also a bad gradient $\mathbf{g}_{\text{bad}}$ with flipped labels. Finally, each malicious client submits the difference $\mathbf{g}_{\text{bad}}-\mathbf{g}_{\text{benign}}$ to the server to achieve the goal that the aggregated global gradient (averaged over all clients) is close to $\mathbf{g}_{\text{bad}}$.
5) DP-LFH. The baseline (Section 3.2) directly combines DP-SGD based momentum with LFH. Each client adds DP noise to the local gradient, and then computes the local momentum to be aggregated with centered clipping by the server.
Appendix F Parameters in Experiments
Basic Parameters.
  • Total number of iterations $T$: 1000 for MNIST and FEMNIST; 2000 for CIFAR-10.
  • Learning rate $\eta_{t}$: For MNIST and FEMNIST datasets, $\eta_{t}$ is linearly reduced from 0.1 to 0.01 w.r.t. iterations. For CIFAR-10 dataset, $\eta_{t}$ is linearly reduced from 0.05 to 0.0025 w.r.t. iterations.
DP-related Parameters.
  • Record-level sampling rate $p_{i}$: 0.05 for all $i$ on MNIST and CIFAR-10; 0.1 for all $i$ on FEMNIST (because each client has fewer data records).
  • Record-level clipping bound $R$: linearly reduced from $R_{0}$ to $0.3R_{0}$ w.r.t. iterations. Note that in Figure 5, the different value of $R$ in x-axis is the value of the above $R_{0}$. For MNIST and FEMNIST, we set $R_{0}=10$ by default. For CIFAR-10, we set $R_{0}=20$ by default, but $R_{0}=15$ only for the case of $\epsilon=2$ in Figure 4.
  • Privacy parameter $\delta$ in DP: $10^{-6}$
  • Noise multiplier $\sigma$: For MNIST (with $T=1000$ and each client has $|\mathcal{D}|_{i}=60000/100=600$ examples), $\sigma\in\{0.15,0.06,0.029,0\}$ for $\epsilon\in\{1,3,8,\text{inf}\}$. For CIFAR-10 (with $T=2000$ and each client has $|\mathcal{D}|_{i}=50000/100=500$ examples), $\sigma\in\{0.14,0.077,0.042,0\}$ for $\epsilon\in\{2,4,9,\text{inf}\}$. For FEMNIST (with $T=1000$ and each client has around $300$ examples), $\sigma\in\{0.16,0.09,0.047,0\}$ for $\epsilon\in\{2,4,9,\text{inf}\}$.
Robustness-related Parameters.
  • Client-level clipping bound $C$ (only for DP-BREM and DP-LFH): linearly reduced from $C_{0}$ to $0.3C_{0}$ w.r.t. iterations, where $C_{0}=1$ for MNIST, $C_{0}=5$ for CIFAR-10, and $C_{0}=2$ for FEMNIST.
  • Momentum parameter $\beta=0.9$ for all three datasets.
Appendix G Useful Lemmas
Lemma 3.
[…]
Proof.
Denote $x_{ki}$ as the $i$-th element of the vector $\bm{x}_{k}$, then
where the inequality is caused by Cauchy-Schwarz inequality. ∎
Lemma 4.
Consider the following optimization problem
$f^{*}=\min_{x_{1},\cdots,x_{K}}~{}~{}\sum\nolimits_{k=1}^{K}c_{k}/x_{k},\qquad\text{s.t.}\quad x_{k}>0,\quad\sum\nolimits_{k=1}^{K}x_{k}=1$
Proof.
Appendix H Proof of Lemma 2 (Aggregation Sensitivity)